OpenStudy (anonymous):

Solve,If two numbers are chosen from (1,2,….25) what is the probability that one of the numbers is twice greater than the other number?

OpenStudy (anonymous):

hmmm... i'm not very good at these but i'll take a stab at it.... the total possible ways of taking 2 numbers out of 25 is 25 choose 2: 25 C 2 now there are 12 ways to choose those two numbers such that one number is twice the other... so my answer(guess) would be: \(\large \frac{12}{\text 25 C 2} \) ????

OpenStudy (anonymous):

The total possible is 25*24 right?

OpenStudy (anonymous):

25 C 2 = (25 !)/[(2!)(23!)] = (25)*(24)/[2] = 25*12

OpenStudy (anonymous):